OpenStudy (anonymous):

lim x -> 1 x^2 - 1/x-1

OpenStudy (lalaly):

x->?

OpenStudy (anonymous):

sorry x -> 1

OpenStudy (lalaly):

\[\lim_{x \rightarrow 1}\frac{(x-1)(x+1)}{x-1}=\lim_{x \rightarrow 1}x+1=2\]

OpenStudy (anonymous):

(x-1)(x+1)/(x-1) cancel the common bracket and put x =1 it will be (x+1)=(1+1)=2

Parth (parthkohli):

Use the L'Hopital by finding derivatives till you get a solution.

OpenStudy (anonymous):

when do we factorize any limit ??

OpenStudy (anonymous):

are there some rules for limits, if anyone can provide .. I will appreciate.